i'm having trouble putting this into an equation, thank you in advance!! i really need help...
In 1962, Wilt Chamberlain set an NBA single-game scoring record. He scored 64 times for a total of 100 points. He made no 3-point shots, but he made several free throws worth 1 point each and several regular shots worth 2 points each. how many free throws did he make? How many 2-point shots did he make?

Answer by ankor@dixie-net.com(22740)

In 1962, Wilt Chamberlain set an NBA single-game scoring record.
He scored 64 times for a total of 100 points. He made no 3-point shots, but he made several free throws worth 1 point each and several regular shots worth 2 points each.
how many free throws did he make?
How many 2-point shots did he make?
:
let x = no. of free throws
let y = no. of two point basket
:
x + y = 64
x = (64-y)
:
1x + 2y = 100
now substitute (64-y) for x, you should be able to solve this now